10) A discrete-time signal is shown in Figure2. Sketch and label carefully the signal x[n]u[3 - n] -3 -2 -1 0 1 2 3 4 5 n
Answer #1


Given signal s -2 3-n20 x6J u [3-n] otterwise {167 otternise 3-n20 u (2-n] otterwise {o otterwise f-n]= u E-n] mu Niplied wi :. xg [3-n] = -2 -1 -1 n=-4 - n= -3 : x6] u[z-n] = りニ-2 75451- り=3